# A Comprehensive Survey of Small Language Models in the Era of Large Language Models: Techniques, Enhancements, Applications, Collaboration with LLMs, and Trustworthiness

### v1

@@ 2. Foundational Concepts in Building Language Models
This section will introduce foundational concepts and background knowledge that are important for language models. We will introduce the basic concepts both in architecture and the training process respectively. The advanced training strategy to improve SLM performance will be introduced in Section 3.
@@ 2.1. Architecture of SLMs
Generally, the architecture of small language models (SLMs) is based on LLMs but optimized for computational efficiency and scalability. SLMs commonly employ the Transformer architecture (Vaswani, 2017) (see Figure 4), which utilizes self-attention mechanisms to manage long-range text dependencies, essential for maintaining performance with constrained resources.
2.1.1. Transformer (Vaswani, 2017) for SLMs
The Transformer’s self-attention mechanism (Vaswani, 2017) allows SLMs to efficiently capture contextual information across longer sequences, even with limited resources. SLM Transformers generally adopt an encoder-decoder structure featuring self-attention mechanisms, feedforward networks, positional embeddings, and layer normalization.
@@ 2.1.1. Transformer (Vaswani, 2017) for SLMs
(1) $\text{MultiHead}({\mathbf{Q}},{\mathbf{K}},{\mathbf{V}})=\text{Concat}(\text{head}_{1},\text{head}_{2},\dots,\text{head}_{h}){\mathbf{W}}^{O},~\text{with}~\text{head}_{i}=\text{Attention}({\mathbf{Q}}{\mathbf{W}}_{i}^{Q},{\mathbf{K}}{\mathbf{W}}_{i}^{K},{\mathbf{V}}{\mathbf{W}}_{i}^{V}),$
@@
Feedforward Network (FFN) comprises two linear transformations separated by a non-linearity, typically modeled as $\text{FFN}(x)=\sigma(x{\mathbf{W}}_{1}+b_{1}){\mathbf{W}}_{2}+b_{2}$. where ${\mathbf{W}}_{1}$ and ${\mathbf{W}}_{2}$ are the weight matrices, and $b_{1}$ and $b_{2}$ are bias terms. $\sigma$ is the activation function, which introduces non-linearity, allowing models to learn complex patterns. Generally, ReLU is used as the activation function. In addition to ReLU, activation functions such as GeLU and SiLU are also used in SLMs to improve performance. We give the details here: (i) ReLU (Rectified Linear Unit) (Agarap, 2018) is defined as $\sigma(x)=\max(0,x)$, which is commonly used for its simplicity and effectiveness. (ii) GELU (Gaussian Error Linear Unit) (Hendrycks and Gimpel, 2016) is defined as $\text{GELU}(x)=x\cdot\Phi(x)=x\cdot\frac{1}{2}\left[1+\text{erf}\left(\frac{x}{\sqrt{2}}\right)\right]$, where $\Phi(x)$ is the standard Gaussian CDF and erf is the error function. It is smoother than ReLU and widely used in models such as BERT (Devlin et al., 2019) and GPT (Radford et al., 2019) for better gradient flow control. Since calculating the Gaussian error function for each neuron is computationally expensive and time consuming, there are approximations using tanh and sigmoid functions, corresponding to $\text{GELU}_{\text{tanh}}$ and SiLU: (iii) GELU with tanh is defined as $\text{GELU}_{\text{tanh}}(x)=0.5\cdot x\cdot\left(1+\tanh\left(\sqrt{\frac{2}{\pi}}\cdot(x+0.044715\cdot x^{3})\right)\right)$. This approximation uses the Tanh function to simplify computations. (iv) SiLU (Sigmoid Linear Unit) (Elfwing et al., 2018) is calculated as $\text{SiLU}(x)=x\cdot\text{sigmoid}(x)=x\cdot\frac{1}{1+e^{-x}}$. It effectively combines the sigmoid function with its input, enhancing modeling capabilities. (v) SwiGLU (Swish-Gated Linear Units) (Shazeer, 2020) integrates the Swish activation function with Gated Linear Units, defined as $\text{SwiGLU}(x)=\text{Swish}(x\cdot W+b)\odot(x\cdot V+c)$ where $W,V$ are the weight matrix and $b,c$ are the bias terms. The Swish function, a smooth non-linear activation, is expressed as $\text{Swish}(x)=x\cdot\text{sigmoid}(x)$. This combination enhances expressiveness and computational efficiency, making it a preferred choice in advanced models such as the Qwen series (Yang et al., 2024e).
Positional Embeddings in Transformer models (Vaswani, 2017) are essential for capturing token order, providing context about relative positions within a sequence. Traditional positional embeddings in the Transformer architecture utilize a sinusoidal function, defined as:
(2) $PE(pos,2i)=\sin\left(\frac{pos}{10000^{2i/d_{\text{model}}}}\right)\quad PE(pos,2i+1)=\cos\left(\frac{pos}{10000^{2i/d_{\text{model}}}}\right)$
@@
where pospos represents the position within the sequence, $i$ is the dimension index, and $d_{\text{model}}$ is the dimensionality of the model. This method alternates between sine and cosine functions across dimensions. BERT-family models (Devlin et al., 2019; Liu et al., 2019), in contrast, employ learned positional embeddings, enhancing flexibility and adaptation to various text structures. To improve the model’s capacity for understanding the relative positions of tokens within a sequence, Rotary Positional Embedding (RoPE) (Su et al., 2024) introduces a rotational matrix to the embeddings. RoPE significantly enhances the positional encoding by maintaining the relative distances through rotational transformations, thus optimizing the model’s interpretative ability regarding sequence dynamics.
Layer Normalization (Lei Ba et al., 2016) stabilizes the training process by normalizing layer outputs, accelerating convergence. Two types of layer normalization are commonly used (Lei Ba et al., 2016): (i) Non-Parametric Layer Norm normalizes inputs using the mean and variance calculated across the layer’s dimensions without learnable parameters as
$\text{LN}(x)=\frac{x-\mu}{\sigma}$
where $\mu$ is the mean and $\sigma$ is the standard deviation of the inputs. Its simplicity makes it ideal for SLMs. (ii) Parametric Layer Norm includes learnable parameters $\gamma$ and $\beta$ for adaptive scaling and bias, enhancing model flexibility:
$\text{PLN}(x)=\gamma\left(\frac{x-\mu}{\sigma}\right)+\beta$
Additionally, RMS Norm (Root Mean Square Layer Normalization) (Zhang and Sennrich, 2019) simplifies the calculation by using the root mean square of inputs, reducing computational demands:
$\text{RMSNorm}(x)=\gamma\frac{x}{\sqrt{\frac{1}{N}\sum_{i=1}^{N}x_{i}^{2}+\epsilon}}+\beta$
where $N$ is the number of inputs, $x_{i}$ is the $i$-th input, and $\epsilon$ is a small constant to prevent division by zero.
2.1.2. Mamba Model (Gu and Dao, 2023; Dao and Gu, 2024)
Mamba leverages a refined version of the Transformer architecture, which includes improvements in multi-head attention and feedforward networks. Specifically, Mamba uses dynamic attention heads, which adjusts the number of active attention heads based on the complexity of the input sequence. This approach mirrors the Multi-Head Attention mechanism within the Transformer but with the added benefit of dynamically scaling the attention resources to match the task’s demands. This dynamic adaptation allows Mamba to maintain high performance while minimizing unnecessary computations for simpler sequences, making it highly efficient. Additionally, Mamba incorporates adaptive feedforward networks, which align closely with the Feedforward Network architecture in Transformers. These adaptive networks dynamically adjust the depth of the feedforward layers based on the input’s complexity, ensuring that computational resources are allocated efficiently. This adaptive mechanism preserves power for more demanding tasks while keeping the model lightweight for simpler tasks, directly improving the efficiency of the Transformer architecture in SLMs.
@@ 2.2. Training SLMs from Scratch
2.2.1. Pretrain
Typically, pre-training paradigms for language models are divided into encoder-based and decoder-based approaches. Encoder-based models, such as BERT (Devlin et al., 2019), utilize Masked Language Modeling (MLM) tasks where the goal is to predict masked tokens within a sentence. This is achieved by maximizing:
[…]
@@ 2.2.1. Pretrain
[…]
where next token is the token that the model aims to predict, context represents the sequence of tokens preceding the token to be predicted, and $\mathbf{h}_{\text{last}}$ is the output from the transformer encoder for the last token in the context. Effective data preprocessing, crucial for optimizing the performance of SLMs trained from scratch, involves meticulous data cleaning and strategic tokenization. Data Cleaning involves techniques such as filtering, deduplication, and noise reduction, which improve data quality and help the model generalize better. Filtering noisy or irrelevant data, addressing outliers, and handling imbalances in the dataset ensure that the training data is both representative and efficient. Deduplication, in particular, helps prevent overfitting by removing repeated instances, making the model more robust with efficient parameter usage. Tokenization Strategies play a vital role in handling diverse vocabularies without increasing model size. Advanced methods such as Byte-Pair Encoding (BPE) (Gage, 1994) and WordPiece (Song et al., 2021) break text into subwords (Devlin et al., 2019), allowing the model to manage rare and compound words efficiently. These strategies ensure that SLMs maintain a balance between vocabulary coverage and model compactness, crucial for improving generalization while minimizing computational demands.
@@ 2.2.2. Fine-Tuning
After the initial training, SLMs are fine-tuned on specific tasks using task-specific data and loss functions. Parameter-efficient fine-tuning methods, such as Low-Rank Adaptation (LoRA), prefix-tuning, and adapter modules, are particularly effective for SLMs. Low-Rank Adaptation (LoRA) (Hu et al., 2021) modifies Transformer weights by introducing trainable low-rank matrices $\mathbf{A}$ and $\mathbf{B}$ for efficient fine-tuning, avoiding significant alterations to pre-trained weights. The update is represented as:
(3) $\Delta\mathbf{W}=\mathbf{A}\mathbf{B}^{\top}$
The fine-tuned weight matrix used in Transformer operations then becomes:
(4) $\mathbf{W}_{\text{ft}}=\mathbf{W}+\alpha\Delta\mathbf{W}$
where $\alpha$ is a scaling factor adjusting the adaptation’s impact, allowing fine-tuning on a smaller set of parameters while retaining the model’s foundational capabilities. Prefix-Tuning (Li and Liang, 2021) prepends learnable prefixes to the input sequence, guiding the model’s attention without altering core model parameters. It is especially useful for generative tasks. Adapter Modules (Houlsby et al., 2019) are small, trainable layers inserted into the pre-trained model. These layers are fine-tuned on task-specific data, allowing the base model to remain fixed while the adapters learn the necessary adjustments. The typical structure of an adapter module includes a down-projection, a non-linearity, and an up-projection:
(5) $\text{Adapter}(\mathbf{h})=\mathbf{h}+\mathbf{W}_{\text{up}}\cdot\sigma(\mathbf{W}_{\text{down}}\cdot\mathbf{h}+\mathbf{b}_{\text{down}})+\mathbf{b}_{\text{up}}$
where $\mathbf{h}$ is the input hidden state, $\mathbf{W}_{\text{down}}$ and $\mathbf{W}_{\text{up}}$ are the projection matrices, $\mathbf{b}_{\text{down}}$ and $\mathbf{b}_{\text{up}}$ are the bias terms, and $\sigma$ is a non-linear activation function.

### v2

@@ 2. Foundational Concepts in Building Language Models
This section will introduce foundational concepts and background knowledge for language models, including the concepts of architecture and the training process, as well as methods for obtaining SLMs from LLMs. The advanced training strategy to improve SLM performance will be introduced in Section 3.
@@ 2.1. Architecture of SLMs
SLMs commonly employ the Transformer architecture (Vaswani, 2017) (see Figure 4), which utilizes self-attention mechanisms to manage long-range text dependencies, essential for maintaining performance with constrained resources. However, due to the attention mechanism, Transformers have large inference cost. Hence, to alleviate the issue, several subquadratic-time architectures such as Mamba (Gu and Dao, 2023), Hymba (Dong et al., 2024), and xLSTM (Beck et al., 2024) are proposed. Next, we will give details of Transformer due to its popularity and briefly introduce newly emerged models.
Figure 4. Transformer architecture (Vaswani, 2017).
2.1.1. Transformer
The Transformer’s self-attention mechanism (Vaswani, 2017) allows language models to efficiently capture contextual information across longer sequences, even with limited resources. The Transformer generally adopts an encoder-decoder structure featuring self-attention mechanisms, feedforward networks, positional embeddings, and layer normalization. The Transformer architecture design tailored for SLMs is detailed in Section 5; this section will provide only foundational concepts.
@@ 2.1.1. Transformer
(1) $\text{MultiHead}({\mathbf{Q}},{\mathbf{K}},{\mathbf{V}})=\text{Concat}(\text{head}_{1},\text{head}_{2},\dots,\text{head}_{h}){\mathbf{W}}^{O},~{}\text{with}~{}\text{head}_{i}=\text{Attention}({\mathbf{Q}}{\mathbf{W}}_{i}^{Q},{\mathbf{K}}{\mathbf{W}}_{i}^{K},{\mathbf{V}}{\mathbf{W}}_{i}^{V})$
@@
Feedforward Network (FFN) comprises two linear transformations separated by a non-linearity, typically modeled as $\text{FFN}(\mathbf{x})=\sigma(\mathbf{x}{\mathbf{W}}_{1}+b_{1}){\mathbf{W}}_{2}+b_{2}$. where ${\mathbf{W}}_{1}$ and ${\mathbf{W}}_{2}$ are the weight matrices, and $b_{1}$ and $b_{2}$ are bias terms. $\sigma$ is the activation function, which introduces non-linearity, allowing models to learn complex patterns. Generally, ReLU is used as the activation function. In addition to ReLU, activation functions such as GeLU and SiLU are also used in SLMs to improve performance. We give the details here: (i) ReLU (Rectified Linear Unit) (Agarap, 2018) is defined as $\sigma(x)=\max(0,x)$, which is commonly used for its simplicity and effectiveness. (ii) GELU (Gaussian Error Linear Unit) (Hendrycks and Gimpel, 2016) is defined as $\text{GELU}(x)=x\cdot\Phi(x)=x\cdot\frac{1}{2}\left[1+\text{erf}\left(\frac{x}{\sqrt{2}}\right)\right]$, where $\Phi(x)$ is the standard Gaussian CDF and erf is the error function. It is smoother than ReLU and widely used in models such as BERT (Devlin et al., 2019) and GPT (Radford et al., 2019) for better gradient flow control. Since calculating the Gaussian error function for each neuron is computationally expensive and time-consuming, there are approximations using tanh and sigmoid functions, corresponding to $\text{GELU}_{\text{tanh}}$ and SiLU: (iii) GELU with tanh is defined as $\text{GELU}_{\text{tanh}}(x)=0.5\cdot x\cdot\left(1+\tanh\left(\sqrt{\frac{2}{\pi}}\cdot(x+0.044715\cdot x^{3})\right)\right)$. This approximation uses the Tanh function to simplify computations. (iv) SiLU (Sigmoid Linear Unit) (Elfwing et al., 2018) is calculated as $\text{SiLU}(x)=x\cdot\text{sigmoid}(x)=x\cdot\frac{1}{1+e^{-x}}$. It effectively combines the sigmoid function with its input, enhancing modeling capabilities. (v) SwiGLU (Swish-Gated Linear Units) (Shazeer, 2020) integrates the Swish activation function with Gated Linear Units, defined as $\text{SwiGLU}(x)=\text{Swish}(x\cdot W+b)\odot(x\cdot V+c)$ where $W,V$ are the weight matrix and $b,c$ are the bias terms. The Swish function is expressed as $\text{Swish}(x)=x\cdot\text{sigmoid}(x)$. This combination enhances expressiveness and computational efficiency, making it a preferred choice in advanced models such as the Qwen series (Yang et al., 2024h).
@@
Positional Embeddings in Transformer models (Vaswani, 2017) are essential for capturing token order, providing context about relative positions within a sequence. Traditional positional embeddings in the Transformer architecture utilize a sinusoidal function, defined as: $PE(pos,2i)=\sin\left(\frac{pos}{10000^{2i/d_{\text{model}}}}\right),\quad PE(pos,2i+1)=\cos\left(\frac{pos}{10000^{2i/d_{\text{model}}}}\right)$ where $pos$ represents the position within the sequence, $i$ is the dimension index, and $d_{\text{model}}$ is the dimensionality of the model. To improve the model’s capacity for understanding the relative positions of tokens within a sequence, Rotary Positional Embedding (RoPE) (Su et al., 2024) introduces a rotational matrix to the embeddings. RoPE significantly enhances the positional encoding by maintaining the relative distances through rotational transformations, thus optimizing the model’s interpretative ability regarding sequence dynamics.
Layer Normalization (Lei Ba et al., 2016) stabilizes the training process by normalizing layer outputs, accelerating convergence. Two types of layer normalization are commonly used (Lei Ba et al., 2016): (i) Non-Parametric Layer Norm normalizes inputs using the mean and variance calculated across the layer’s dimensions without learnable parameters as $\text{LN}(x)=\frac{x-\mu}{\sigma}$ where $\mu$ is the mean and $\sigma$ is the standard deviation of the inputs. Its simplicity makes it ideal for SLMs. (ii) Parametric Layer Norm includes learnable parameters $\gamma$ and $\beta$ for adaptive scaling and bias, enhancing model flexibility: $\text{PLN}(x)=\gamma\left(\frac{x-\mu}{\sigma}\right)+\beta$ Additionally, RMS Norm (Root Mean Square Layer Normalization) (Zhang and Sennrich, 2019) simplifies the calculation by using the root mean square of inputs, reducing computational demands: $\text{RMSNorm}(x)=\gamma\frac{x}{\sqrt{\frac{1}{N}\sum_{i=1}^{N}x_{i}^{2}+\epsilon}}+\beta$ where $N$ is the number of inputs, $x_{i}$ is the $i$-th input, and $\epsilon$ is a small constant to prevent division by zero.
Figure 5. Mamba 1 architecture (Gu and Dao, 2023).
2.1.2. Mamba
The attention mechanism in Transformer suffers from a drawback: it requires recalculating attention scores with every previous token for each new token generated during inference, leading to quadratic time complexity. This increases the inference cost as sequence lengths grow. In contrast, Mamba (Gu and Dao, 2023; Dao and Gu, 2024), based on state space models (SSMs) (Kalman, 1960) which are a superclass of recurrent neural networks (RNNs), rely only on the last hidden state for generating the next token, enabling faster inference speeds, as shown in Figure 5. To address the Linear Time Invariant nature of traditional SSMs, which hinders their ability to focus on or ignore specific inputs, Mamba improves SSMs with a dynamic selection mechanism. This mechanism selectively filters out irrelevant information while retaining essential data, tailored to the content of the input. Leveraging this selective SSM foundation, Mamba adeptly captures complex global relationships within sequence data. Due to its focus on the immediate previous hidden state, as opposed to Transformer which requires access to all previous hidden states, Mamba achieves a higher utilization rate of model parameters. This makes it more suitable for SLMs. However, we identify two drawbacks of Mamba: (i) its focus on selectively capturing global information may compromise performance on tasks that require nuanced understanding, such as detailed sentiment analysis or complex entity recognition; and (ii) to balance inference speed, Mamba’s recurrent structure primarily encodes static global information, which limits its effectiveness in handling multi-round tasks within a single query, such as interactive dialogue systems or iterative problem-solving scenarios.
In language modeling, Mamba 1 (Gu and Dao, 2023) is pre-trained on the Pile dataset (Gao et al., 2020) using the training recipe from (Brown et al., 2020) and ranges from 125M to 1.3B parameters. It outperforms comparable models such as Pythia (Biderman et al., 2023) and RWKV (Peng et al., 2023a) in various tasks; for instance, Mamba-1.4B achieves a 32.8% accuracy on the Arc-Challenge (Clark et al., 2018) dataset, surpassing Pythia-1.4B’s 28.5% and RWKV-1.5B’s 29.4%. Mamba 2 (Dao and Gu, 2024) develops a theoretical framework linking SSMs with attention mechanisms through structured semi-separable matrices, enhancing the selective SSM to achieve 2-8x faster speeds while competing with Transformer models. Training and configuration for Mamba 2 align with Mamba 1. Additionally, Mamba-series models are applied widely across different fields (Zuo et al., 2024; Qu et al., 2024a; Jiang et al., 2023; Qu et al., 2024b). Other follow-up Mamba-based language models such as Falcon Mamba 7B (Zuo et al., 2024) and Mistral 7B (Jiang et al., 2023) also demonstrate the efficiency of the architecture for NLP tasks. Falcon Mamba 7B scales Mamba’s long-sequence processing capabilities to extensive language data, reducing memory overhead and making it ideal for long-form NLP tasks. Similarly, Mistral 7B incorporates Mamba’s efficient sequence handling ability to improve processing speed and computational efficiency in long-context NLP tasks, showcasing Mamba’s scalability and practicality for large-scale language modeling.
Figure 6. Hymba (Dong et al., 2024) architecture.
2.1.3. Hymba
Attention heads in the Transformer facilitate high-resolution recall, while SSM heads in Mamba efficiently summarize context. To balance performance and efficiency for SLMs, Hymba (Dong et al., 2024) integrates both attention and SSM heads within the same layer, allowing for parallel and complementary processing of inputs, as depicted in Figure 6. This hybrid-head approach enables each layer to simultaneously leverage the high-resolution recall of attention heads and the contextual summarization of SSMs, increasing the model’s flexibility and expressiveness in managing diverse information flows and memory access patterns.
Hymba has been developed in models of varying sizes—125M, 350M, and 1.5B, trained on a combination of the DCLM-Baseline-1.0 (Li et al., 2024b), SmolLM-Corpus (Ben Allal et al., 2024), and a proprietary high-quality dataset, with token counts of 1 trillion, 250 billion, and 50 billion, respectively. The models incorporate the Warmup-Stable-Decay (WSD) learning rate scheduler (Hu et al., 2024b) and the data annealing technique (Dubey et al., 2024) to ensure stable pretraining, conducted on 128 NVIDIA A100 GPUs. The 1.5B base model was post-trained using full finetuning (FFT), followed by direct preference optimization (DPO) (Rafailov et al., 2024) to develop the Hymba-1.5B-Instruct model. In commonsense reasoning tasks, the Hymba 1.5B model surpasses Llama-3.2-3B (AI, 2024) by achieving 1.32% higher average accuracy, requiring an 11.67$\times$ smaller cache size, and delivering a 3.49$\times$ increase in processing speed.
2.1.4. xLSTM
Long Short-Term Memory (LSTM) (Hochreiter and Schmidhuber, 1996) shares a conceptual similarity with Mamba that achieves success in language modeling through the introduction of time-dependent weights. This similarity raises an intriguing question: how effective would LSTMs be at language modeling if scaled to billions of parameters, incorporating advanced techniques from modern LLMs while addressing known LSTM limitations? Inspired by this question, Beck et al. (Beck et al., 2024) propose xLSTM architecture, which performs favorably compared to state-of-the-art Transformers and SSMs in empirical evaluations. To address the limitations of LSTM, xLSTM designs exponential gates to enhance effectiveness with long sequences, expand memory cells from scalars to matrices to increase storage capacity, and remove memory mixing to enable parallel processing.
To test the language modeling capabilities of xLSTM scaled to billions of parameters, it is trained on a large dataset comprising 300 billion tokens from SlimPajama (Soboleva et al., 2023) across various model sizes (125M, 350M, 760M, 1.3B). The performance of pre-trained xLSTM is compared against RWKV-4 (Peng et al., 2023a), Llama (Inan et al., 2023), and Mamba (Gu and Dao, 2023) across 571 text domains of the PALOMA benchmark (Magnusson et al., 2023) and various downstream tasks. Across all model sizes and the majority of tasks, xLSTM consistently outperforms the others, suggesting that larger xLSTM models could become formidable competitors to existing Large Language Models that utilize Transformer technology.
@@ 2.2. Training SLMs from Scratch
2.2.1. Pre-training
Typically, the pre-training paradigm for language models is divided into encoder-based and decoder-based approaches. Encoder-based models, such as BERT (Devlin et al., 2019), utilize Masked Language Modeling (MLM) tasks where the goal is to predict masked tokens within a sentence. This is achieved by maximizing:
$P(\text{masked token}\mid\text{context})=\text{softmax}(\mathbf{W}\cdot\mathbf{h}_{\text{mask}}+b),$
@@ 2.2.1. Pre-training
$P(\text{next token}\mid\text{context})=\text{softmax}(\mathbf{W}\cdot\mathbf{h}_{\text{last}}+b),$
where next token is the token that the model aims to predict, context represents the sequence of tokens preceding the token to be predicted, and $\mathbf{h}_{\text{last}}$ is the output from the transformer encoder for the last token in the context. Effective data preprocessing, crucial for optimizing the performance of SLMs trained from scratch, involves meticulous data cleaning and strategic tokenization.
Data Cleaning involves techniques such as filtering, deduplication, and noise reduction, which improve data quality and help the model generalize better. Filtering noisy or irrelevant data, addressing outliers, and handling imbalances in the dataset ensure that the training data is both representative and efficient. Deduplication, in particular, helps prevent overfitting by removing repeated instances, making the model more robust with efficient parameter usage.
Tokenization plays a vital role in handling diverse vocabularies without increasing model size. Advanced methods such as Byte-Pair Encoding (BPE) (Gage, 1994) and WordPiece (Song et al., 2021) break text into subwords (Devlin et al., 2019), allowing the model to manage rare and compound words efficiently. These strategies ensure that SLMs maintain a balance between vocabulary coverage and model compactness, crucial for improving generalization while minimizing computational demands.
@@ 2.2.2. Fine-Tuning
After the initial training, SLMs are fine-tuned on specific tasks using task-specific data and loss functions. Parameter-efficient fine-tuning methods (Hu et al., 2021; Li and Liang, 2021; Houlsby et al., 2019; Gu et al., 2024b), such as Low-Rank Adaptation (LoRA), prefix-tuning, and adapter modules, are particularly effective for SLMs. Low-Rank Adaptation (LoRA) (Hu et al., 2021) modifies Transformer weights by introducing trainable low-rank matrices $\mathbf{A}$ and $\mathbf{B}$ for efficient fine-tuning, avoiding significant alterations to pre-trained weights. The update is represented as: $\Delta\mathbf{W}=\mathbf{A}\mathbf{B}^{\top}$ The fine-tuned weight matrix used in Transformer operations then becomes: $\mathbf{W}_{\text{ft}}=\mathbf{W}+\alpha\Delta\mathbf{W}$ where $\alpha$ is a scaling factor adjusting the adaptation’s impact, allowing fine-tuning on a smaller set of parameters while retaining the model’s foundational capabilities. Prefix-Tuning (Li and Liang, 2021) prepends learnable prefixes to the input sequence, guiding the model’s attention without altering core model parameters. It is especially useful for generative tasks. Adapter Modules (Houlsby et al., 2019) are small, trainable layers inserted into the pre-trained model. These layers are fine-tuned on task-specific data, allowing the base model to remain fixed while the adapters learn the necessary adjustments. The typical structure of an adapter module includes a down-projection, a non-linearity, and an up-projection: $\text{Adapter}(\mathbf{h})=\mathbf{h}+\mathbf{W}_{\text{up}}\cdot\sigma(\mathbf{W}_{\text{down}}\cdot\mathbf{h}+\mathbf{b}_{\text{down}})+\mathbf{b}_{\text{up}}$ where $\mathbf{h}$ is the input hidden state, $\mathbf{W}_{\text{down}}$ and $\mathbf{W}_{\text{up}}$ are the projection matrices, $\mathbf{b}_{\text{down}}$ and $\mathbf{b}_{\text{up}}$ are the bias terms, and $\sigma$ is a non-linear activation function.